In Dallas, some fire trucks were painted yellow (instead of red) to heighten their visibility. During a test period, the fleet of red fire trucks made 153,348 runs and had 20 accidents, while the fleet of yellow fire trucks made 135,035 runs and had 4 accidents. At α = .01, did the yellow fire trucks have a significantly lower accident rate? (a) State the hypotheses. (b) State the decision rule and sketch it. (c) Find the sample proportions and z test statistic. (d) Make a decision. (e) Find the p-value and interpret it. (f ) If statistically significant, do you think the difference is large enough to be important? If so, to whom, and why? (g) Is the normality assumption fulfilled? Explain.
Statistic Red Fire Trucks Yellow Fire Trucks
Number of accidents x1 = 20 accidents x2 = 4 accidents
Number of fire runs n1 = 153,348 runs n2 = 135,035 runs

Answer by stanbon(75887)

In Dallas, some fire trucks were painted yellow (instead of red) to heighten their visibility. During a test period, the fleet of red fire trucks made 153,348 runs and had 20 accidents, while the fleet of yellow fire trucks made 135,035 runs and had 4 accidents. At α = .01, did the yellow fire trucks have a significantly lower accident rate?
(a) State the hypotheses.
Ho: p(red)-p(yellow) = 0
Ha: p(red)-p(yellow) > 0
-----------------------------------------------
(b) State the decision rule and sketch it.
Reject Ho if the test statistic is greater than z = 2.3263
-----------------------------------------------
(c) Find the sample proportions and z test statistic.
p-hat(red) = 20/153348 ; p-hat(yellow) = 4/135,035
z-test statistic: 2.961
----------------------------------
(d) Make a decision.
Reject Ho
-------------------------------
(e) Find the p-value and interpret it.
p-value: P(z>2.961) = 0.0015
Only 0.15% of test results could provide stronger evidence for
rejecting Ho.
----------------------------------
(f ) If statistically significant, do you think the difference is large enough to be important? If so, to whom, and why?
Yes, it supports painting all engines yellow
----------------------------------------------------
(g) Is the normality assumption fulfilled? Explain.
The sample sizes are large enough to support the conclusion.
-------------------------------------------------------
Cheers,
Stan H.
-------------------------------------------------------
Statistic Red Fire Trucks Yellow Fire Trucks
Number of accidents x1 = 20 accidents x2 = 4 accidents
Number of fire runs n1 = 153,348 runs n2 = 135,035 runs